# Interpretable Measurement of CNN Deep Feature Density using Copula and the Generalized Characteristic Function

### v1

@@ 5 Conclusion
We present an empirical analysis of the density distribution of deep CCN features using a novel non-parametric copula analysis technique that models the GCF by measuring the orthogonal moments. We show that our approach is able to model the marginal and interdependence terms of feature density after each major Conv+ReLU block of resnet18, resnet50, and vgg19. Moreover, we do so without any unjustified or unverified parametric assumptions as to the shape of the marginal or interdependence terms. We report empirical findings on the observed marginal distributions and copula density interdependence terms as a function of network depth. In our analysis of marginals, we observe that features after major Conv+ReLU blocks exhibit both zero and non-zero features. Furthermore, we demonstrate through hypothesis testing that the non-zero features for the deeper layers of the network more closely fit the Weibull or Gamma distribution versus the Gaussian or Uniform distribution. Weibull and Gamma distributions generalize the Exponential distribution which also significantly outperforms the Gaussian and Uniform distributions in the deeper layers. Our analysis of the copula interdependence shows that pairs of features in the deeper layers of the network also exhibit an unusual form of statistical dependence, for which these features are highly independent throughout their typical value ranges, yet become strongly dependent (either correlated or anti-correlated) for extremely large feature values. We observe that the Legendre and Fourier GCF techniques are able to better model the copula interdependence density relative to the histogram technique for groups of four features with resnet18 and imagenette2 for the deeper layers of the network. Our approach is the first purely empirical technique to model the multivariate probability density distribution of deep CNN features. We believe that empirical analysis of the feature density distribution will lead to a better understanding of CNN feature representations. Moreover empirical feature density has the potential to lead to a new branch of generative methods that model the full joint distribution of the native CNN feature space, thereby enabling off-the-shelf CNN architectures to attain all of the benefits of generative techniques.
@@ 6 Discussion
Moreover, the uncorrelated nature of typical value features, yet strong correlation of extreme value features, would suggest that future work should revisit the assumption that the typical set is in fact the most descriptive set of features for a given image, as modeling of feature density has overwhelmingly emphasized evaluation of the density distribution of typical valued features. We believe that modeling of the distribution of extreme valued features may also be highly important. Moreover, given the assumption of an exponential distribution, the removal of outlier features under the assumption of a Multi-Variate Gaussian has the potential to accidentally eliminate strong computer vision signals of foreground target detection which would be an unwanted side effect.
@@ 7 Acknowledgment
We would like to thank the University of Miami Frost Institute for Data Science and Computing for their support of this research. We would also like to thank Prof. Vanessa Aguilar-Pulido, Odelia Schwartz, Stefan Wuchty, and Dilip Sarkar, as well as students Anthony Givans, Nicolas Solages, and Wenqian Zhu who have all contributed to this work through their encouragement and valuable feedback.

### v2

@@ 5 Conclusion
We present an empirical analysis of the density distribution of deep CNN features through direct measurement of the GCF using a novel non-parametric approach that combines copula analysis with the MOM. We demonstrate that our approach is able to model the marginal and interdependence terms of feature density after each major Conv+ReLU block of resnet18, resnet50, and vgg19. Moreover, as a non-parametric technique, we do not introduce overly-restrictive assumptions as to the shape of the marginal or interdependence terms. We report empirical findings on the observed marginal distributions and copula density interdependence terms as a function of network depth. In our analysis of marginals, we observe that features after major Conv+ReLU blocks exhibit both zero and non-zero features. Furthermore, we demonstrate through hypothesis testing that the non-zero features for the deeper layers of the network more closely fit the Weibull or gamma distribution versus the Gaussian or uniform distribution. Weibull and gamma distributions generalize the exponential distribution which also significantly outperforms the Gaussian and uniform distributions in the deeper layers. Our analysis of the copula interdependence shows that pairs of features in the deeper layers of the network also exhibit an unusual form of statistical dependence, for which these features are highly independent throughout their typical value ranges, yet become strongly dependent (either correlated or anti-correlated) for extremely large feature values. We observe that the Legendre and Fourier GCF techniques are able to better model the copula interdependence density relative to the histogram technique for groups of four features with resnet18 and imagenette2 for the deeper layers of the network. Our approach is the first purely empirical technique to model the multivariate probability density distribution of deep CNN features. We believe that empirical analysis of the feature density distribution will lead to a better understanding of CNN feature representations. Moreover empirical feature density has the potential to lead to a new branch of generative methods that model the full joint distribution of the native CNN feature space, thereby enabling off-the-shelf CNN architectures to attain all of the benefits of generative techniques.
@@ 6 Discussion
Moreover, the uncorrelated nature of typical value features, yet strong correlation of extreme value features, would suggest that future work should revisit the assumption that the typical set is in fact the most descriptive set of features for a given image, as modeling of feature density has overwhelmingly emphasized evaluation of the density distribution of typical valued features. We believe that modeling of the distribution of extreme valued features may also be highly important. Moreover, given the assumption of an exponential distribution, the removal of outlier features under the assumption of an MVG has the potential to accidentally eliminate strong computer vision signals of foreground target detection which would be an unwanted side effect.
@@ 7 Acknowledgment
We would like to thank the University of Miami Frost Institute for Data Science and Computing for their support of this research. We would also like to thank Prof. Vanessa Aguilar-Pulido, Odelia Schwartz, Stefan Wuchty, and Dilip Sarkar, as well as students Anthony Givans, Nicholas Solages, and Wenqian Zhu who have all contributed to this work through their encouragement and valuable feedback.